OpenStudy (anonymous):

y = f(x), where the roots are 1 and 3, and the minimum point is (2,-1). sketch the graph, marking the points of intersection of the axes. so i sketched it, but don't know how to find the y intercept, and therefore the formula :P

OpenStudy (anonymous):

is this a quadratic function? degree 2 right?

OpenStudy (anonymous):

you can easily find the equation. in fact you have too much information. method 1: vertex is \((2,-1)\) tells you it looks like \[f(x)=a(x-2)^2-1\] the only thing you don't know it \(a\) but you know \(f(3)=0\) so \[a(3-2)^2-1=0\] \[a-1=0\] \[a=1\] and your function is \[f(x)=(x-2)^2-1\]

OpenStudy (anonymous):

method 2: you have the zeros are 1 and 3 so it must look like \[f(x)=a(x-1)(x-3)\] and again all you are missing is \(a\) this time use the fact that \(f(2)=-1\) so get \[f(2)=a(2-1)(2-3)=-1\] \[-a=-1\] \[a=0\] as before, so it is \[f(x)=(x-1)(x-3)\] expand either one to put in standard form

OpenStudy (anonymous):

if any step is not clear let me know

OpenStudy (anonymous):

well, in method 2 you meant to say -a = -1, therefore a = 1, not 0 :P but anyhoo, thanks, i understood it really well :)